OpenStudy (anonymous):

Evaluate the indefinite integral. integral of 6x/x^4+1 dx Please help me! Thanks!

sam (.sam.):

u-sub u=x^2 du=2xdx \[3\int\limits \frac{1}{u^2+1} \, du\]

sam (.sam.):

trig substitution

sam (.sam.):

u=tan(y) \[y=\tan^{-1}u\] Integrate \[3 \tan ^{-1}(u)+c\] \[3 \tan ^{-1}\left(x^2\right)+c\]

OpenStudy (anonymous):

Hmm.. I see.. That's how you do it! Thanks so much! I've been doing this for days and I couldn't get it.. I see where I got confuse.. Thanks so much! :D

sam (.sam.):

yw